OpenStudy (anonymous):

Find the equation of the line of best fit from the data in the table. Round the slope and the y-intercept to the nearest hundredth. x 12,14,16,18,20 y54,53,55,54,56

OpenStudy (anonymous):

A line contains at least two points. 2 A plane contains at least three noncollinear points. Which option best classifies Saul’s statements? Statement 1 is a theorem because it can be proved and Statement 2 is a postulate because it is a true fact. Statement 1 and Statement 2 are theorems because they can be proved. Statement 1 and Statement 2 are postulates because they are true facts. Statement 1 is a postulate because it is a true fact and Statement 2 is a theorem because it can be proved. --------------------------------------------------------------------------------

OpenStudy (anonymous):

Let \[ y = m x + b\\ m = \frac { 5 \sum_{i=1}^5 x_i y_i- \left( \sum_{i=1}^5 x_i \right) \left( \sum_{i=1}^5 y_i \right) } {5 \left( \sum_{i=1}^5 x_i^2 \right)- \left( \sum_{i=1}^5 x_i \right)^2}=1.15\\ b= \frac {\left( \sum_{i=1}^5 y_i \right)- m \left( \sum_{i=1}^5 x_i \right) } { 5 }=37.8\\ y= 1.15 x + 37.8 \]